OpenStudy (anonymous):

If h(x) = 2x + 1, find the value of x if h(x) = 11. 5 23 6 10

OpenStudy (anonymous):

\[2x+1=11\] solve for \(x\) in two steps

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

well first subtract 1 from both sides then divide by 2. so what is 11 -1 ?

OpenStudy (anonymous):

11-1=10 10/2=5 -_-